OpenStudy (anonymous):

Which of the following is not = to 0 a. cos(pi/2) b. tan(0) c. cos(pi/4)-sin(pi/4) d. sin^-1(pi) e. cos^-1(1) I know the answer is either d or e

OpenStudy (anonymous):

sin pi is 0 cos 1 is (pi/2))

OpenStudy (dumbcow):

sorry that was not what they wanted... cos(?) = 1 sin(?) = pi

OpenStudy (anonymous):

2pi 0

OpenStudy (dumbcow):

no sin(0) = 0

OpenStudy (dumbcow):

cos(0) = 1

OpenStudy (dumbcow):

answer then must be D

OpenStudy (anonymous):

but it's inverse sin and inverse cos the inverse is confusing to me

OpenStudy (dumbcow):

ok the output of the inverse is the angle \[\sin \theta = x \rightarrow \theta = \sin^{-1} x\] so when they ask about an inverse, think sin of what angle will give me that number example \[\cos^{-1} \frac{\sqrt{2}}{2}\] rewrite it as \[\cos \theta = \frac{\sqrt{2}}{2}\] then from the unit circle you know the angle must be pi/4 \[\cos^{-1} \frac{\sqrt{2}}{2} = \frac{\pi}{4}\]

OpenStudy (anonymous):

so because it's cos^-1(1) the angle is 0 therefore cos^-1(1)=0?

OpenStudy (dumbcow):

correct

OpenStudy (anonymous):

thanks!

OpenStudy (dumbcow):

yw :)